Find the values for m and b in the following system so that the solution to the system is (-3, 4).
5x + 7y = b
mx + y = 22
Ok, I need some help here. This one has me confussed and frustrated at the same time. Thanks I appreciate any help you could give me.

Answer by Earlsdon(6294)

Find the values of m and b in the following given that a solution is (-3, 4)
5x+7y = b Substitute the x- and y-coordinates of the givenpoint (-3, 4) and solve for b.
5(-3) + 7(4) = b Simplify.
-15 + 28 = b
13 = b
mx + y = 22 As above, substitute the x- and y-coordinates of the given point and solve for m.
m(-3) + 4 = 22
-3m + 4 = 22 Subtract 4 from both sides.
-3m = 18 Divide both sides by -3.
m = -6